### v1

# Model Editing with Graph-Based External Memory

Abstract
Large language models (LLMs) have revolutionized natural language processing, yet their practical utility is often limited by persistent issues of hallucinations and outdated parametric knowledge. Although post-training model editing offers a pathway for dynamic updates, existing methods frequently suffer from overfitting and catastrophic forgetting. To tackle these challenges, we propose a novel framework that leverages hyperbolic geometry and graph neural networks for precise and stable model edits. We introduce HYPE [^1] HYPE (HYperbolic Parameter Editing), which comprises three key components: (i) Hyperbolic Graph Construction, which uses Poincaré embeddings to represent knowledge triples in hyperbolic space, preserving hierarchical relationships and preventing unintended side effects by ensuring that edits to parent concepts do not inadvertently affect child concepts; (ii) Möbius-Transformed Updates, which apply hyperbolic addition to propagate edits while maintaining structural consistency within the hyperbolic manifold, unlike conventional Euclidean updates that distort relational distances; and (iii) Dual Stabilization, which combines gradient masking and periodic GNN parameter resetting to prevent catastrophic forgetting by focusing updates on critical parameters and preserving long-term knowledge. Experiments on CounterFact, CounterFact+, and MQuAKE with GPT-J and GPT2-XL demonstrate that HYPE significantly enhances edit stability, factual accuracy, and multi-hop reasoning.
Model Editing with Graph-Based External Memory
Yash Kumar Atri University of Virginia atri@virginia.edu                        Ahmed Alaa UC Berkeley & UCSF amalaa@berkeley.edu                        Tom Hartvigsen University of Virginia hartvigsen@virginia.edu
1 Introduction
Large language models (LLMs) such as GPT-3 Brown et al. (2020), PaLM Chowdhery et al. (2023), and LLaMA Touvron et al. (2023) have revolutionized natural language processing (NLP), enabling unprecedented capabilities in text generation Li et al. (2024a), reasoning Chowdhery et al. (2023), and contextual understanding Zhao et al. (2024). These models underpin a wide range of applications, from conversational agents Liu et al. (2024) to knowledge-intensive tasks like question answering Phukan et al. (2024); Shah et al. (2024); Gao et al. (2024) and summarization Atri et al. (2023a, 2021, b, c); Dey et al. (2020); Atri et al. (2023d). However, their reliance on static, pre-trained parametric knowledge renders them prone to generating factual inaccuracies Wang et al. (2024), hallucinations Ji et al. (2023), and outdated information—critical limitations in real-world deployments where accuracy and timeliness are paramount Lazaridou et al. (2021); Atri et al. (2025). While fine-tuning on updated data can mitigate these issues, the computational and data demands of retraining billion-parameter models Kaplan et al. (2020) render this approach impractical for real-time knowledge updates. Instead, post-training model editing has emerged as a promising alternative, enabling targeted modifications to a model’s parametric knowledge without full retraining Mitchell et al. (2022a).
Existing model editing methods, such as ROME Meng et al. (2022a) and MEMIT Meng et al. (2022b), rely on Euclidean geometry to modify specific facts. While these methods achieve localized edits, they struggle with hierarchical relationships due to Euclidean space’s flat structure. For example, ROME enforces memorization via equality constraints, while MEMIT uses least-squares optimization, both of which fail to preserve relational depth Yang et al. (2024). This leads to geometric mismatch, where edits distort semantically related knowledge Nickel and Kiela (2017). Additionally, Euclidean updates in overparameterized models like GPT-2 Radford et al. (2019a) cause update instability, propagating unintended changes through dense connections Kaplan et al. (2020). Finally, these methods suffer from contextual fragility, lacking mechanisms to maintain global coherence during edits Zhong et al. (2023).
Hyperbolic geometry, in contrast, naturally encodes hierarchical relationships through its exponential growth property, making it well-suited for modeling linguistic knowledge Nickel and Kiela (2017). Unlike Euclidean spaces, where distances are linear, hyperbolic space expands exponentially, allowing child nodes to be placed far from parent nodes while preserving their relational proximity Ganea et al. (2018). This property enables precise parameter updates that do not propagate unintended side effects. For instance, modifying a parent concept in hyperbolic space affects only its immediate vicinity, preserving the stability of its hierarchical descendants. Furthermore, hyperbolic operations like Möbius addition Ungar (2013) ensure updates remain on the manifold, avoiding the geometric distortion caused by Euclidean vector addition.
Motivated by these insights and the inherent advantages of hyperbolic geometry, we propose HYPE (HYperbolic Parameter Editing), a novel framework that leverages hyperbolic geometry Ganea et al. (2018) for model editing through three core components. First, Hyperbolic Graph Construction projects knowledge graph into a learnable curvature space using Poincaré embeddings Nickel and Kiela (2017), preserving hierarchical relationships—such as parent-child and whole-part associations—via exponential mapping and hyperbolic distances. Second, Möbius-Transformed Ungar (2013) Updates adjust model parameters using hyperbolic addition, a geometric operation that preserves the hierarchical structure of the data. Unlike Euclidean vector addition, which can distort relationships in hierarchical data, Möbius addition ensures that edits remain within the hyperbolic manifold. This prevents unintended side effects by maintaining the relative distances between related concepts. Third, Dual Stabilization combines gradient-based sparsification Frankle and Carbin (2019), which masks negligible gradient updates, with periodic resetting of graph network parameters to prevent the overwriting of critical pre-existing knowledge during the integration of new edits. By unifying hyperbolic embeddings, Möbius-Transformed updates, and robust stabilization techniques, HYPE achieves precise and resilient model edits.
We summarize our contributions as follows:
  1. 1.
    HYPE is the first approach to integrate hyperbolic geometry with graph-based model editing, leveraging Poincaré embeddings to project knowledge graphs into a learnable curvature space that preserves complex hierarchical relationships during model updates.
  2. 2.
    HYPE introduces a novel editing mechanism that applies Möbius transformations for parameter updates, ensuring that edits remain consistent with the hyperbolic manifold and preserving the inherent hierarchical structure—thereby preventing the distortions common with Euclidean vector addition.
  3. 3.
    We validate HYPE across three widely benchmarked datasets—CounterFact Meng et al. (2022a), CounterFact+ Yao et al. (2023b), and MQuAKE Zhong et al. (2023)—and two popular LLMs—GPT-J Wang and Komatsuzaki (2021), and GPT2-XL Radford et al. (2019b)—demonstrating superior factual accuracy, parameter efficiency, and edit stability compared to existing model editing methods.
Code available at: https://github.com/yashkumaratri/HYPE
2 Related Work
Post-Training Model Editing: Model editing techniques broadly fall into parameter-modifying and parameter-preserving approaches. Parameter-modifying methods, such as locate-then-edit strategies Meng et al. (2022a, b); Hartvigsen et al. (2023); Kolbeinsson et al. (2025) and meta-learning frameworks Mitchell et al. (2022a); Cao et al. (2021), adjust model parameters via hypernetworks or rank-one weight updates. Zhang et al. (2024) enhances specificity using knowledge graphs, but these methods typically treat parameter changes as isolated scalar operations (Ma et al., 2024; Gu et al., 2024), disregarding structured knowledge representations in LLMs (Cohen et al., 2023). This often leads to unintended side effects (Yao et al., 2023a; Kolbeinsson et al., 2025) and poor support for compositional updates (Hase and Bansal, 2023).
Parameter-preserving approaches avoid direct weight updates by leveraging external memories or prompts Mitchell et al. (2022b); Huang et al. (2023); Madaan et al. (2022); Zheng et al. (2023), while continual learning methods Zeng et al. (2019); Farajtabar et al. (2020) use orthogonal projections to mitigate catastrophic forgetting. However, these strategies require architectural modifications or fail to maintain long-term coherence in factual updates.
Figure 1: The illustration delineates our proposed model, HYPE. We begin by constructing a hyperbolic knowledge graph (b) using Poincaré embeddings to encode hierarchical relationships. When an edit is required, we apply Möbius transformations (c) to update the weights while ensuring curvature-aware consistency. To maintain stability, Dual stabilization strategy (d) removes transient or spurious updates. The edited knowledge is then integrated into the model (e), preserving factual accuracy and structural integrity.
Graph-Based Model Editing: Knowledge graphs provide a structured approach to model editing by aligning factual updates with relational dependencies. Cohen et al. (2023) aligns LLM attention with Wikidata triples, while Cao et al. (2021) edits GNNs through subgraph swaps. Zhang et al. (2024) integrates knowledge graphs into transformer layers but relies on Euclidean embeddings, distorting hierarchical relationships. Similarly, Qin et al. (2023) models higher-order dependencies via hypergraphs but lacks geometric constraints to maintain relational consistency.
Hyperbolic Model Editing: Hyperbolic embeddings have proven effective for encoding hierarchical knowledge Chami et al. (2019), particularly in NLP Valentino et al. (2024) and knowledge graph applications Chami et al. (2020). Chen et al. (2022) introduces hyperbolic attention for LLMs requiring full retraining, while Cohen et al. (2023) links LLMs to graphs using Euclidean projections, which fails to capture hyperbolic curvature.
HYPE is the first framework to unify hyperbolic embeddings and Möbius updates for model editing. Unlike Euclidean methods (e.g., ROME Meng et al. (2022b), MEMIT Meng et al. (2022b), GLAME Zhang et al. (2024)), which struggle with hierarchical edits and stability, HYPE leverages hyperbolic geometry to preserve relational consistency. Compared to parameter-preserving approaches, it avoids architectural overhead while maintaining edit precision. Experiments show HYPE achieves +9.12 higher Edit Quality Score (EDS) and +3.59 better efficacy than state-of-the-art methods (Table 1), demonstrating the benefits of curvature-aware updates.
3 Methodology
In this section, we detail the methodology underlying HYPE, our proposed framework for hyperbolic model editing.
3.1 Hyperbolic Graph Construction
We first initialize a knowledge graph using Wikidata Vrandečić and Krötzsch (2014) triples, which consist of subject-relation-object tuples $\{(s,r,o)\ ...\}$. These triples encode factual knowledge, where $s$ represents an entity (e.g., "Albert Einstein"), $o$ represents another entity or concept associated with $s$ (e.g., "Theory of Relativity"), and $r$ defines the semantic relationship between them (e.g., "contributed to"). To construct meaningful representations, we first obtain Euclidean embeddings using a graph neural network (GNN), capturing relational and structural dependencies. For any input string $x$ (which may represent an entity or a relation), let $\mathbf{v}_{\text{euc}}(x)$ denote its Euclidean embedding. We then project these embeddings into hyperbolic space via the exponential map on the Poincaré ball Nickel and Kiela (2017)
$\mathbf{v}_{\text{hyp}}(x)$ $=\exp^{c}_{0}\left(\mathbf{v}_{\text{euc}}(x)\right)$ (1)
$=\tanh\left(\sqrt{c}\|\mathbf{v}_{\text{euc}}(x)\|\right)\cdot\frac{\mathbf{v}_{\text{euc}}(x)}{\sqrt{c}\|\mathbf{v}_{\text{euc}}(x)\|}$
where $c$ is the curvature parameter (set to $1.0$ in our experiments). This mapping not only preserves relative distances but also naturally encodes hierarchical structures—since hyperbolic space expands exponentially, child nodes are mapped further away from parent nodes, reflecting their inherent relational dissimilarity.
For relation embeddings, we extract a set of unique relations $\{r_{1},r_{2},\dots,r_{m}\}$ from the triples. Each relation $r$ is first embedded in Euclidean space to yield $\mathbf{r}_{\text{euc}}(r)$ and then projected to hyperbolic space as:
$\mathbf{r}_{\text{hyp}}(r)=\exp^{c}_{0}\left(\mathbf{r}_{\text{euc}}(r)\right).$ (2)
These hyperbolic relation embeddings serve as edge features in our graph.
We then construct a directed graph $G=(V,E)$, where each unique entity encountered in the Wikipedia triples is assigned a node $v\in V$ with an associated feature $\phi(v)$ defined as:
$\phi(v)=\exp^{c}_{0}\left(\mathbf{v}_{\text{euc}}(v)\right).$ (3)
For every triple $(s,r,o)$, an edge $(s,o)\in E$ is created. The edge feature corresponding to this edge is given by the hyperbolic relation embedding:
$\mathbf{e}_{s,o}=\exp^{c}_{0}\left(\mathbf{r}_{\text{euc}}(r)\right),$ (4)
and a relation-to-index mapping is maintained to assign a type ID to each relation.
To stabilize subsequent graph neural network (GNN) computations, self-loops are added to each node, and the in-degree of each node is computed to generate a normalization factor:
$\text{norm}(v)=\text{deg}(v)^{-1},$ (5)
which is incorporated into the node features. Furthermore, we apply a persistence filter to retain topologically significant embeddings:
[…]
where $\tau$ is a persistence threshold (learnable parameter), and $\sigma(\cdot)$ is the sigmoid function. The filter is applied to the initial relation embeddings:
$\mathbf{a}=\text{persistent\_filter}(\mathbf{r}_{\text{hyp}}).$ (7)
This ensures that only structurally important relations contribute to the model update process.
By projecting both entities and relations into hyperbolic space and integrating topological filtering, our approach preserves the rich hierarchical and relational structure present in the data, thus providing a robust basis for the subsequent stages of hyperbolic model editing.
3.2 Möbius-Transformed Weight Update
Traditional weight updates of the form $\mathbf{w}^{\prime}=\mathbf{w}+\Delta$ fail to respect the curvature of hyperbolic space. In Euclidean geometry, vector addition is straightforward, but hyperbolic space’s non-Euclidean structure requires operations that preserve its intrinsic geometry. To address this, we employ Möbius addition, which ensures updates remain on the hyperbolic manifold while maintaining hierarchical relationships encoded in the data. Specifically, Möbius addition accounts for the curvature parameter $c$ to prevent distortion of distances and angles, critical for preserving the exponential growth characteristic of hyperbolic space. This operation guarantees that the updated weight vector $\mathbf{w}_{\text{new}}$ adheres to the manifold’s constraints, enabling stable and context-aware edits. In a space with curvature $c$, the Möbius addition of $\mathbf{w}$ and $\Delta$ is defined as
$\mathbf{w}_{\text{new}}$ $=\mathbf{w}\oplus^{c}\Delta$ (8)
$=\frac{(1+2c\langle\mathbf{w},\Delta\rangle+c\|\Delta\|^{2})\mathbf{w}}{1+2c\langle\mathbf{w},\Delta\rangle+c^{2}\|\mathbf{w}\|^{2}\|\Delta\|^{2}}$
$\quad+\frac{(1-c\|\mathbf{w}\|^{2})\Delta}{1+2c\langle\mathbf{w},\Delta\rangle+c^{2}\|\mathbf{w}\|^{2}\|\Delta\|^{2}}$
[…]
The update term $\Delta$ is computed using the rank-1 algorithm Meng et al. (2022a). First, we derive two vectors: a left update vector $\mathbf{u}\in\mathbb{R}^{m}$ and a right update vector $\mathbf{v}\in\mathbb{R}^{n}$. Their outer product forms the base update:
$\Delta_{0}=\mathbf{u}\otimes\mathbf{v},\quad\text{with }(\mathbf{u}\otimes\mathbf{v})_{ij}=u_{i}\,v_{j}.$ (9)
This base update is scaled by a residual factor $\gamma$, resulting in
$\Delta_{1}=\gamma\,\Delta_{0}.$ (10)
To ensure that only significant gradients contribute to the update, we compute for each output feature the average gradient magnitude:
$g_{i}=\frac{1}{n}\sum_{j=1}^{n}\Big{|}\nabla_{w_{ij}}\mathcal{L}_{\text{edit}}\Big{|},$
and define a soft gradient mask via the sigmoid function:
$m_{i}=\sigma\Big{(}g_{i}-\tau_{g}\Big{)},\quad\text{with }\sigma(x)=\frac{1}{1+e^{-x}},$ (11)
where $\tau_{g}$ is a predefined gradient persistence threshold. By broadcasting the mask $\mathbf{m}$ across the corresponding dimensions, the final update matrix is given by
$\Delta=\Delta_{1}\odot\mathbf{m},$ (12)
where $\odot$ denotes element-wise multiplication. Substituting Eq. (12) into Eq. (8) yields the final weight update:
$\mathbf{w}_{\text{new}}=\mathbf{w}\oplus^{c}\Delta.$ (13)
This Möbius-transformed update procedure ensures that the modifications to $\mathbf{w}$ respect the underlying hyperbolic geometry, thereby preserving the hierarchical structure encoded in the model.
3.3 Dual Stabilization Strategy
To prevent catastrophic forgetting and maintain the geometric consistency of the model, we introduce a dual stabilization strategy comprising hyperbolic projection and periodic resetting of the graph neural network (GNN) parameters.
Hyperbolic Projection:
After applying the Möbius update, it is crucial to ensure that the updated weights remain within the valid region of the Poincaré ball. We achieve this by projecting the updated weights onto the Poincaré ball $\mathbb{D}_{c}^{d}$. The projection operator is defined as
$\text{proj}(\mathbf{w})=\min\left\{1,\frac{1/\sqrt{c}}{\|\mathbf{w}\|}\right\}\mathbf{w},$ (14)
which guarantees that $\|\mathbf{w}\|\leq 1/\sqrt{c}$. Thus, the stabilized weight is given by
$\mathbf{w}_{\text{final}}=\text{proj}\Bigl{(}\mathbf{w}\oplus^{c}\Delta\Bigr{)}.$ (15)
Graph Neural Network Reset: The update directions $\mathbf{u}$ and $\mathbf{v}$ are computed using a GNN operating on the hyperbolic graph constructed from Wikipedia triples. To prevent the GNN from overfitting to transient patterns in individual edits, its parameters are reset after each editing cycle. Formally, if $\theta_{\text{gnn}}^{(0)}$ denotes the initial GNN parameters, then after each update step $t$, we enforce
$\theta_{\text{gnn}}^{(t+1)}=\theta_{\text{gnn}}^{(0)}.$ (16)
This resetting mechanism prevents the architecture from overfitting to individual edits, ensuring that transient adaptations do not accumulate and disrupt the model’s internal coherence.
Method Model Counterfact/+ MQuAKE
Eff Gen Spec Port+ EDS 2-hops 3-hops 4-hops Avg
Zeroshot GPT-J 15.47 17.43 80.96 10.27 37.95 14.67 22.19 10.43 15.76
[…]
MEND 44.76 44.83 52.28 12.68 47.29 13.86 11.24 9.62 11.57
ROME 55.77 52.57 50.49 28.43 52.94 32.63 29.04 17.33 26.33
MEMIT 95.59 92.64 61.73 28.84 83.32 35.47 26.93 15.38 25.93
PMET 83.57 84.24 52.25 27.61 73.35 31.47 24.67 13.21 23.12
RAE 94.84 84.02 70.05 29.68 83.30 32.53 26.08 14.92 24.51
HYPE 99.43 98.35 79.47 29.83 92.42 46.68 39.73 23.13 36.51
$\Delta$(HYPE - best base) $\uparrow 3.84$ $\uparrow 5.71$ $\uparrow 9.42$ $\uparrow 0.15$ $\uparrow 9.10$ $\uparrow 11.21$ $\uparrow 10.69$ $\uparrow 5.80$ $\uparrow 10.18$
Zeroshot GPT2-XL 21.56 23.61 76.17 10.06 40.45 23.61 21.95 14.37 19.98
[…]
MEND 58.65 53.26 47.73 14.08 53.21 26.49 24.71 14.93 22.04
ROME 98.63 94.73 73.53 21.29 88.96 38.52 30.28 16.74 28.51
MEMIT 92.84 78.48 76.33 18.63 82.13 34.62 26.08 15.88 25.53
PMET 91.66 77.08 75.19 17.33 81.69 32.17 23.68 13.83 23.23
RAE 89.34 76.41 63.18 23.19 76.31 30.42 25.27 14.31 23.33
HYPE 99.57 97.18 77.86 24.53 91.54 43.62 33.83 22.79 33.41
$\Delta$(HYPE - best base) $\uparrow 0.94$ $\uparrow 2.45$ $\uparrow 1.53$ $\uparrow 1.34$ $\uparrow 2.58$ $\uparrow 5.10$ $\uparrow 3.55$ $\uparrow 6.05$ $\uparrow 4.90$
Table 1: Comprehensive evaluation of HYPE against model editing baselines on GPT-J and GPT-2 XL. We assess performance across three benchmarks: (1) Counterfact – measuring Efficacy (Eff), Generalization (Gen), Specificity (Spec), and Edit Quality Score (EDS); (2) Counterfact+ – evaluating Portability (Port+); and (3) MQuAKE – testing multi-hop reasoning across 2-, 3-, and 4-hop tasks. Results indicate that HYPE consistently outperforms all baselines, with particularly strong gains in multi-hop reasoning. Higher values denote better performance.
4 Datasets, Baselines and Evaluation
We evaluate HYPE on three widely benchmarked datasets – 1) CounterFact Meng et al. (2022a), which assesses factual accuracy, specificity, and generalization; 2) CounterFact+ Yao et al. (2023b), which evaluates edit portability across paraphrased queries; 3) MQuAKE Zhong et al. (2023), which tests multi-hop reasoning capabilities. We compare HYPE against seven state-of-the-art baselines – (i) Zeroshot (unmodified model), (ii) FT (Fine-Tuning), (iii) MEND Mitchell et al. (2022a), (iv) ROME Meng et al. (2022a), (v) MEMIT Meng et al. (2022b), (vi) PMET, and (vii) RAE Shi et al. (2024). Further details on datasets and baselines are provided in Appendix A.
We measure HYPE’s performance using the following metrics – 1) Efficacy (Eff): Quantifies the accuracy of factual edits by measuring the model’s ability to correctly answer questions after updates. 2) Generalization (Gen): Assesses robustness across paraphrased queries and reasoning tasks, ensuring edits apply consistently to varied inputs. 3) Specificity (Spec): Measures unintended alterations to unrelated knowledge, evaluating the model’s ability to avoid cascading errors. 4) Portability (Port+): Evaluates the model’s ability to transfer edits across different contexts, ensuring updates remain effective under rephrased questions (used for benchmarking CounterFact+). 5) Edit Quality Score (EDS): A composite metric defined as the harmonic mean of efficacy, generalization, and specificity, providing a holistic view of edit performance. 6) Multi-Hop Efficacy: For MQuAKE, efficacy is measured over 2-hop, 3-hop, and 4-hop reasoning tasks, assessing the model’s ability to handle complex, multi-step queries.
5 Experimental Results
This section presents the evaluation of HYPE against seven baselines—Zeroshot, FT, MEND Mitchell et al. (2022a), ROME Meng et al. (2022a), MEMIT Meng et al. (2022b), PMET Li et al. (2024b), and RAE Shi et al. (2024) — on three benchmark datasets: (1) CounterFact Meng et al. (2022a), assessed using Efficacy (Eff), Generalization (Gen), Specificity (Spec), and Edit Quality Score (EDS); (2) CounterFact+ Yao et al. (2023b), evaluated with Portability Score (Port+); and (3) MQuAKE Zhong et al. (2023), measured using Multi-Hop Efficacy (2,3,4-hop). We test HYPE with two LLMs — GPT-J Wang and Komatsuzaki (2021) and GPT-2XL Radford et al. (2019a). The results (Table 1) demonstrate that HYPE consistently outperforms all baselines across factual accuracy, edit stability, and multi-hop reasoning tasks. We also assess the individual contribution of each proposed module in the ablation study (Table 2 and Section 5.4). Further details on experimentation setup and evaluation protocols are provided in Appendix B.
Method Model Counterfact/+ MQuAKE
Eff Gen Spec Port+ EDS 2-hops 3-hops 4-hops Avg
HYPE 99.43 98.35 79.47 29.83 92.42 46.68 39.73 23.13 36.51
HYPE w/o DualS GPT-J 97.45 96.12 77.34 24.98 90.30 44.32 38.21 21.45 34.66
HYPE w/o HGraph 96.23 95.45 76.54 24.56 89.41 43.45 37.32 20.89 33.89
HYPE w/o Mobius updates 92.89 90.45 69.56 23.08 84.30 40.12 35.12 19.23 31.49
HYPE w/o HGraph & Mobius 90.12 88.34 68.56 22.34 82.34 38.45 33.45 18.76 30.22
HYPE w/o HGraph & Mobius & DualS 88.45 86.34 66.54 21.34 80.44 36.45 31.45 16.78 28.23
HYPE 99.57 97.18 77.86 24.53 91.54 43.62 33.83 22.79 33.41
HYPE w/o DualS GPT2-XL 93.45 91.34 73.45 23.45 86.08 42.34 32.34 21.34 32.01
HYPE w/o HGraph 92.34 90.45 72.34 23.00 85.04 41.34 31.34 20.34 31.01
HYPE w/o Mobius updates 88.45 86.34 68.54 22.34 81.11 39.45 30.45 19.45 29.78
HYPE w/o HGraph & Mobius 86.45 84.34 67.45 21.34 79.41 38.45 29.45 18.45 28.78
HYPE w/o HGraph & Mobius & DualS 84.45 82.34 66.45 20.34 77.75 37.45 28.45 17.45 27.78
Table 2: Ablation study on GPT-J and GPT2-XL architectures. The highlighted rows show full HYPE performance; removing components (Dual Stabilization (DualS), Hyperbolic Graph (HGraph), and Mobius updates) progressively degrades performance.
5.1 Counterfact and Counterfact+ Results
On the Counterfact dataset, HYPE achieves an efficacy score of 99.43 and generalization score of 98.35 when evaluated on the GPT-J model. These results represent a +3.84 improvement in efficacy and a +5.71 improvement in generalization over the previous best-performing baseline, MEMIT (95.59 Eff, 92.64 Gen). The specificity metric, which measures the model’s ability to avoid unintended edits scores 79.47, a +9.42 points improvement over MEMIT’s 70.05. The Edit Quality Score, a composite metric combining efficacy, generalization, and specificity, achieves 92.42, outperforming MEMIT’s 83.32 by +9.10 points. For Counterfact+, which evaluates the portability of edits across different contexts, HYPE attains a score of 29.83, surpassing MEMIT’s 28.84 by +0.15. These results highlight HYPE’s ability to maintain factual accuracy and coherence while minimizing unintended side effects.
When evaluated on the GPT2-XL model, HYPE continues to demonstrate superior performance. With an efficacy score of 99.57, a generalization score of 97.18, and a specificity of 77.86, it outperforms the best baselines by +0.94 (efficacy over ROME), +2.45 (generalization over ROME), and +1.53 (specificity over MEMIT). The Edit Quality Score reaches 91.54, a +2.58 improvement over ROME’s 88.96. These results indicate that HYPE maintains its effectiveness across different model architectures and scales, as shown in Table 1.
5.2 MQuAKE Results
The MQuAKE dataset evaluates multi-hop reasoning capabilities, requiring models to answer complex questions that involve multiple steps of inference. HYPE achieves significant improvements across all multi-hop tasks. For 2-hop reasoning, HYPE attains an efficacy score of 46.68 on GPT-J, outperforming MEMIT’s 35.47 by +11.21 points. On 3-hop reasoning, HYPE scores 39.73, a +10.69 score improvement over ROME’s 29.04. For 4-hop reasoning, HYPE achieves 23.13, surpassing ROME’s 17.33 by +5.80 improvement score. The average accuracy across all multi-hop tasks reaches 36.51, a +10.18 improvement over ROME’s efficacy of 26.33. These results highlight HYPE’s ability to preserve hierarchical relationships and relational consistency during edits, which are critical for multi-hop reasoning.
On the GPT2-XL model, HYPE achieves 43.62, 33.83, and 22.79 efficacy score in 2, 3, 4-hop, respectively. These results represent improvements of +5.10, +3.55, and +6.05 over ROME’s 38.52, 30.28, and 16.74 efficacy scores respectively. The average accuracy of 33.41 represents a +4.90 improvement over ROME’s 28.51. These results further demonstrate HYPE’s robustness across different model architectures and its ability to maintain coherence in complex reasoning tasks.
5.3 Model-Specific Performance
HYPE demonstrates consistent improvements across both GPT-J and GPT2-XL models. On GPT-J, the Edit Quality Score increases by +9.10 over MEMIT (92.42 vs. 83.32), with efficacy improving by +3.84 and generalization by +5.71 points. On GPT2-XL, the EDS improves by +2.58 over ROME (91.54 vs. 88.96), with efficacy increasing by +0.94 (99.57 vs. 98.63) and generalization by +2.45 (97.18 vs. 94.73). These results indicate that HYPE’s improvements are not limited to a specific model architecture but generalize across different scales and configurations. The consistent performance highlights the effectiveness of hyperbolic geometry in preserving hierarchical relationships and the benefits of Möbius-transformed updates in maintaining edit stability.
5.4 Analysis of Ablation Results
The ablation study results in Table 2 demonstrate the critical role of each component in HYPE’s architecture. On the GPT-J model, removing Dual Stabilization (HYPE w/o DualS) reduces efficacy by 1.98 points and specificity by 2.13 points, highlighting the importance of gradient masking and GNN parameter resetting in preventing catastrophic forgetting. Further ablating the Hyperbolic Graph (HYPE w/o HGraph) causes a -3.20 drop in efficacy and a -2.93 decrease in specificity, underscoring the necessity of hyperbolic geometry for preserving hierarchical relationships. When both the Hyperbolic Graph and Möbius updates are removed, performance declines sharply, with efficacy falling by -9.31 and specificity by -10.93. The most drastic degradation occurs when all three components are removed (HYPE w/o HGraph & Mobius & DualS), resulting in a -10.98 points drop in efficacy and a -12.93 decrease in specificity. Similar trends are observed on the larger GPT2-XL model, where removing Dual Stabilization reduces efficacy by -6.12 and specificity by -4.41, while ablating the Hyperbolic Graph and Möbius updates causes an -11.12 decline in efficacy and a -9.41 reduction in specificity. The full ablation (HYPE w/o HGraph & Mobius & DualS) leads to a -15.12 decrease in efficacy and a -11.41 drop in specificity. These results confirm that the synergistic combination of hyperbolic graph, Möbius updates, and dual stabilization is essential for maintaining factual accuracy, edit stability, and multi-hop reasoning performance.
6 Discussion and Analysis
HYPE demonstrates strong performance on localized edits while maintaining edit stability, even in cases where multi-hop reasoning is challenging. Below, we analyze representative examples to illustrate the model’s capabilities. The complete model outputs for the analyzed cases are presented in Appendix Section C
6.1 Qualitative Analysis of Successful Edits
[…]
Figure 2: Left: EDS peaks at $c=1.0$ due to an optimal balance between expansion and numerical stability. Right: Edit success rate declines for $\tau>0.5$ due to underfitting.
6.2 Qualitative Analysis of Challenging Edits
Although HYPE encounters challenges in multi-hop reasoning in some samples, it consistently demonstrates impressive edit stability and specificity. For instance, in Case ID 560 ("British Railways Board"), the model successfully updated the headquarters from London to Prague, achieving 100% correctness in rewrite prompts. While paraphrase and neighborhood edits showed some inconsistency, the dual stabilization module effectively mitigated catastrophic forgetting by preserving critical knowledge. Similarly, in Case ID 264 ("Toyota RAV4"), the model succeeded in 60% of neighborhood edits, illustrating its partial propagation capabilities. These results highlight HYPE’s ability to handle complex edits while maintaining stability, a strength notably absent in Euclidean approaches.
6.3 Impact of Hyperbolic Geometry
1) Curvature Sensitivity: Curvature $c$ defines the rate of expansion in hyperbolic space. Experiments with $c\in\{0.5,1.0,2.0\}$ (Figure 2, left) showed optimal results at $c=1.0$, where Edit Quality Score (EDS) reached 92.4 points. Higher curvature ($c=2.0$) caused gradient instability, reducing EDS to 80.1, while lower curvature ($c=0.5$) resulted in insufficient relational separation (EDS = 85.2).
2) Persistence Filtering: This method uses a persistence threshold $\tau$ to filter out topologically insignificant updates. The persistence threshold $\tau$ in Eq. (6) critically affected edit stability (Figure 2, right). Lower $\tau$ values ($<0.5$) led to overfitting, with EDS dropping to 82.4 at $\tau=0.8$. Optimal performance (92.4 EDS) is achieved at $\tau=0.5$, balancing specificity and generality. Higher thresholds ($>0.8$) caused underfitting by filtering out critical updates.
7 Conclusion
In this paper, we proposed HYPE, a novel post-training model editing framework that leverages hyperbolic geometry and graph neural networks to deliver precise and stable updates to large language models. Our approach overcomes critical issues in existing methods, including overfitting and catastrophic forgetting, by constructing a hyperbolic graph with Poincaré embeddings, employing a Möbius-transformed update strategy, and integrating a dual stabilization mechanism with gradient masking and GNN parameter resetting. Evaluations on CounterFact, CounterFact+, and MQuAKE with GPT-J and GPT2-XL show that HYPE significantly enhances factual accuracy and parameter efficiency while mitigating unintended side effects. Future work will explore domain-specific constraints where hyperbolic editing could enforce critical factual accuracy.
8 Limitations
HYPE faces several limitations. While it improves edit stability and factual accuracy, the computational overhead of hyperbolic operations and persistence filtering is higher than Euclidean methods, although this is offset by reduced retraining needs. Current implementation focuses on medium-sized LLMs (e.g., GPT-J, GPT2-XL); scaling to larger models (e.g., GPT-4, LLaMA-3-70b) would require distributed optimization and increased memory resources. The framework relies on structured knowledge graphs (e.g., Wikipedia triples), which may not capture implicit or evolving knowledge, and assumes the availability of clean, curated triples. Additionally, while effective for hierarchical edits, fine-grained edits to individual tokens or low-level linguistic features may require additional mechanisms beyond hyperbolic graph updates.
9 Ethics Statement
The benchmarked datasets – Counterfact, Counterfact+, and MQuAKE are available under open-source licenses. The baselines are also covered by the open-source licenses. All experiments are performed on one Nvidia A6000 (48GB) GPU.
Appendix
Appendix A Datasets and Baselines
In this appendix, we provide detailed descriptions of the datasets and baselines used in our experiments and the metrics employed to evaluate the performance of our proposed model, HYPE.
A.1 Datasets
We conduct our experiments on three widely used datasets for model editing: CounterFact, CounterFact+, and MQuAKE.
  • CounterFact dataset Meng et al. (2022a) contains over 3,000 instances designed to evaluate a model’s ability to perform accurate, specific, and generalizable factual edits. It assesses efficacy (Eff), generalization (Gen), specificity (Spec), and edit quality score (EDS) by measuring a model’s ability to update facts without unintended side effects.
  • CounterFact+ dataset Yao et al. (2023b) extends CounterFact by evaluating edit portability across paraphrased queries. It contains over 1,000 paraphrased questions and uses the portability (Port+) metric to measure a model’s ability to transfer edits across different linguistic formulations.
  • MQuAKE dataset Zhong et al. (2023) tests multi-hop reasoning with over 3,000 complex, multi-step questions (2-hop, 3-hop, 4-hop). Each instance includes one or more edits and associated multi-hop questions, requiring models to leverage edited knowledge to answer multi-step queries. The dataset uses the multi-hop efficacy metric to evaluate performance, with 2-hop, 3-hop, and 4-hop questions.
A.2 Baselines
We compare HYPE against seven state-of-the-art model editing techniques:
  1. 1.
    Zeroshot: The unmodified model, serving as a baseline for unedited performance.
  2. 2.
    FT (Fine-Tuning): The full model fine-tuned on edited data, updating all parameters.
  3. 3.
    MEND Mitchell et al. (2022a): Updates gradient-based weights to align edits with factual knowledge while preserving other information.
  4. 4.
    ROME Meng et al. (2022a): Modifies specific relation parameters in the model, assuming that knowledge can be localized to a single layer.
  5. 5.
    MEMIT Meng et al. (2022b): Performs memory-injected editing, updating weights across multiple layers to incorporate new knowledge.
  6. 6.
    PMET Li et al. (2024b): Parameter-efficient editing using low-rank weight updates to minimize computational overhead.
  7. 7.
    RAE Shi et al. (2024): Uses a Retrieval-Augmented Generation (RAG) based approach, incorporating external knowledge during inference.
A.3 Evaluation Metrics
We define the following metrics to assess model editing performance:
Efficacy (Eff)
Measures the accuracy of factual edits by evaluating whether the model answers updated questions correctly:
$\mathbb{E}_{i}\Bigl{[}\,\mathbb{P}\bigl{(}f_{\theta}(o_{i}\mid(s_{i},r_{i}))\bigr{)}>\mathbb{P}\bigl{(}f_{\theta}(o_{c}^{i}\mid(s_{i},r_{i}))\bigr{)}\Bigr{]}$
Generalization (Gen)
Assesses how consistently the model applies edits to paraphrased queries:
$\mathbb{E}_{i}\Bigl{[}\,\mathbb{P}\bigl{(}f_{\theta}(o_{i}\mid N(s_{i},r_{i}))\bigr{)}>\mathbb{P}\bigl{(}f_{\theta}(o_{c}^{i}\mid N(s_{i},r_{i}))\bigr{)}\Bigr{]}$
Specificity (Spec)
Quantifies unintended alterations to unrelated knowledge:
$\mathbb{E}_{i}\Bigl{[}\,\mathbb{P}\bigl{(}f_{\theta}(o_{i}^{c}\mid O(s_{i},r_{i}))\bigr{)}>\mathbb{P}\bigl{(}f_{\theta}(o_{c}^{i}\mid O(s_{i},r_{i}))\bigr{)}\Bigr{]}$
Edit Quality Score (EDS)
A composite measure given by the harmonic mean of Efficacy, Generalization, and Specificity.
Portability (Port+)
Evaluates how well edits transfer across rephrased contexts:
$\mathbb{E}_{i}\Bigl{[}\,\mathbb{P}\bigl{(}f_{\theta}(o_{i}\mid N(s_{i},r_{i}))\bigr{)}>\mathbb{P}\bigl{(}f_{\theta}(o_{c}^{i}\mid N(s_{i},r_{i}))\bigr{)}\Bigr{]}$
We use the similar Efficacy metric to compute the n-hop Efficacy scores over the MQuAKE dataset.
Appendix B Experimental Setup and Evaluation Protocol
B.1 Implementation Details
The experiments are conducted using PyTorch 2.0, DGL 1.1, and HuggingFace Transformers 4.30 on an NVIDIA A6000 GPU with 48GB memory. We evaluate HYPE on three datasets: CounterFact, CounterFact+, and MQuAKE, using two base models: GPT-J and GPT2-XL. The hyperparameters for each configuration were carefully tuned to optimize performance on each dataset.
Parameter CF CF+ MQuAKE
Layers 5 9 5/9
GNN Grad Steps 25–35 35–50 25–50
GNN Loss Layer 27/47 47 27/47
Learning Rate (gnn_lr) $15e-$ 1 $15e-$ 1 $15e-$ 1
Weight Decay $11e-$ $15e-$ 1 $15e-$ 1 $11e-$ $15e-$ 1
Dropout (Attn/Feat) 0.2/0.2–0.4 0.2/0.3–0.4 0.2/0.3–0.4
KL Factor 0.0625–0.075 0.0625–0.075 0.0725–0.075
Early Stopping Loss $23e-$ $24e-$ 2 $24e-$ 2/$5e-3$ $23e-$ 2/$5e-3$
Table 3: Hyperparameters for CounterFact (CF), CounterFact+ (CF+), and MQuAKE for GPT-J/GPT2-XL models.
For the hyperbolic settings, we use the Poincaré Ball model from the geoopt [^2] https://github.com/geoopt/geoopt library with a curvature parameter c=1.0. Hyperbolic operations (Möbius addition) are applied during the model updates. The hyperbolic space is initialized with a learnable curvature, allowing the model to adapt to the hierarchical structure of the data.
The hyperparameters for each dataset and model configuration are detailed in Table 3. The number of GNN gradient steps and GNN loss layer are adjusted based on the dataset and model complexity. The learning rate, weight decay, and dropout rates are also tuned to achieve optimal performance. Early stopping is implemented using the ablated loss thresholds to prevent overfitting.
B.2 Evaluation Protocol
For the evaluation benchmarks, we preserve the dataset splits proposed in the original works for CounterFact Meng et al. (2022a) and MQuAKE Zhong et al. (2023). For CounterFact, we evaluate our method on the first 7500 records for both GPT-J and GPT2-XL models. For CounterFact+, we utilize the 1031 samples provided for testing. For MQuAKE, we follow the settings used in Zhong et al. (2023) and use a subset of 3000 entries. These entries are evenly distributed across 2-hop, 3-hop, and 4-hop questions, with each category comprising 1000 entries. This distribution ensures a balanced evaluation of the model’s performance across different levels of reasoning complexity.
B.3 Hyperbolic Graph Construction
The foundation of our model lies in constructing a hyperbolic graph that accurately represents hierarchical relationships in the data. We initialize a Poincaré Ball model with a curvature parameter $c=1.0$, enabling effective embedding of knowledge triples in hyperbolic space. This curvature allows the model to capture hierarchical structures more effectively than Euclidean spaces, as distances in hyperbolic space grow exponentially, aligning well with tree-like structures in knowledge graphs.
To process the knowledge triples (following (Zhang et al., 2024)), we embed each entity (subject and object) into the hyperbolic space using the exponential map:
$\mathbf{v}_{\text{hyp}}=\exp^{c}_{0}(\mathbf{v}_{\text{eucl}})=\tanh\left(\sqrt{c}|\mathbf{v}_{\text{eucl}}|\right)\cdot\frac{\mathbf{v}_{\text{eucl}}}{\sqrt{c}|\mathbf{v}_{\text{eucl}}|}$ (17)
where $\mathbf{v}_{\text{eucl}}$ is the Euclidean entity vector, and $\mathbf{v}_{\text{hyp}}$ is its hyperbolic counterpart. This transformation preserves hierarchical relationships in the embedding space. Similarly, relation embeddings are projected onto the Poincaré Ball using the same exponential map, serving as edge features to encode directional dependencies.
With node and edge features generated, the graph is constructed by defining nodes for each unique entity and adding edges based on relation types, while incorporating self-loops to enhance message passing. Node features are normalized as:
$\mathbf{n}_{i}=\frac{1}{\sqrt{|\mathcal{N}(i)|}}\sum_{j\in\mathcal{N}(i)}\mathbf{h}_{j}$ (18)
where $\mathbf{n}_{i}$ is the normalized feature for node $i$, $\mathcal{N}(i)$ denotes its neighbors, and $\mathbf{h}_{j}$ is the feature of node $j$, preventing gradient instability.
To refine the graph, we apply a persistence-based filtering mechanism that prunes weak edges based on geometric significance in hyperbolic space, reducing noise and enhancing structure. Edge features are normalized for consistent scaling, ensuring balance across relation types, and the resulting hyperbolic graph effectively captures hierarchical and semantic structures.
Appendix C Sample Model Outputs
We present four model outputs from the CounterFact dataset, generated using HYPE.
Listing 1: Output from model for sample 983
"case_id": 983,
"grouped_case_ids": [983],
"num_edits": 1,
"requested_rewrite": {
"prompt": "{} plays the instrument",
"relation_id": "P1303",
"target_new": {"str": "violin", "id": "Q8355"},
"target_true": {"str": "guitar", "id": "Q6607"},
"subject": "Larry Knechtel"
"time": 10.52,
"post": {
"rewrite_prompts_probs": [{"target_new": 0.0192, "target_true": 6.16}],
"paraphrase_prompts_probs": [{"target_new": 1.13, "target_true": 3.08}, {"target_new": 4.43, "target_true": 11.02}],
"neighborhood_prompts_probs": [{"target_new": 8.69, "target_true": 7.15}]
Listing 2: Output from model for sample 729
"case_id": 729,
"grouped_case_ids": [729],
"num_edits": 1,
"requested_rewrite": {
"prompt": "{} works as",
"relation_id": "P106",
"target_new": {"str": "astronomer", "id": "Q11063"},
"target_true": {"str": "poet", "id": "Q49757"},
"subject": "Johann von Rist"
"time": 10.57,
"post": {
"rewrite_prompts_probs": [{"target_new": 0.0136, "target_true": 14.64}],
"paraphrase_prompts_probs": [{"target_new": 2.08, "target_true": 12.86}],
"neighborhood_prompts_probs": [{"target_new": 12.24, "target_true": 11.59}]
Listing 3: Output from model for sample 560
"case_id": 560,
"grouped_case_ids": [560],
"num_edits": 1,
"requested_rewrite": {
"prompt": "{}’s headquarters are in",
"relation_id": "P159",
"target_new": {"str": "Prague", "id": "Q1085"},
"target_true": {"str": "London", "id": "Q84"},
"subject": "British Railways Board"
"time": 44.47,
"post": {
"rewrite_prompts_probs": [{"target_new": 8.77e-05, "target_true": 12.56}],
"paraphrase_prompts_probs": [{"target_new": 0.0009, "target_true": 9.14}],
"neighborhood_prompts_probs": [{"target_new": 8.06, "target_true": 2.31}]
Listing 4: Output from model for sample 264
"case_id": 264,
"grouped_case_ids": [264],
"num_edits": 1,
"requested_rewrite": {
"prompt": "{} is created by",
"relation_id": "P176",
"target_new": {"str": "Volvo", "id": "Q215293"},
"target_true": {"str": "Toyota", "id": "Q53268"},
"subject": "Toyota RAV4"
"time": 38.36,
"post": {
"rewrite_prompts_probs": [{"target_new": 0.0001, "target_true": 13.30}],
"paraphrase_prompts_probs": [{"target_new": 0.0002, "target_true": 11.79}],
"neighborhood_prompts_probs": [{"target_new": 8.07, "target_true": 1.95}]
Algorithm HYPE Algorithm
1:procedure HyperbolicGraph($\{(s,r,o)\}$)
2:     Pretrained Euclidean embeddings: $\mathbf{v}_{\text{euc}}(x)$ for entities and relations
3:     for each entity $x$ do
4:         Compute hyperbolic embedding: $\mathbf{v}_{\text{hyp}}(x)=\exp_{0}^{c}(\mathbf{v}_{\text{euc}}(x))$
5:     end for
6:     for each relation $r$ do
7:         Project to hyperbolic space: $\mathbf{r}_{\text{hyp}}(r)=\exp_{0}^{c}(\mathbf{r}_{\text{euc}}(r))$
8:     end for
9:     Construct graph $G=(V,E)$ with edges and relation features
10:     Add self-loops, compute $\text{norm}(v)=\text{deg}(v)^{-1}$
11:     Apply persistence filter: $\mathbf{a}=\sigma(\|\mathbf{r}_{\text{hyp}}\|-\tau)$
12:end procedure
13:procedure MöbiusUpdate($\mathbf{w},\mathcal{L}$)
14:     Compute gradient magnitude $\mathbf{g}$ and mask $\mathbf{m}=\sigma(\mathbf{g}-\tau_{g})$
15:     Compute update vectors $\mathbf{u},\mathbf{v}$ and $\Delta=\gamma(\mathbf{u}\otimes\mathbf{v})\odot\mathbf{m}$
16:     Update weights: $\mathbf{w}_{\text{new}}=\mathbf{w}\oplus^{c}\Delta$
17:end procedure
18:procedure Stabilization($\mathbf{w}$, $\theta_{\text{gnn}}$)
19:     Project weights: $\mathbf{w}_{\text{final}}=\text{proj}(\mathbf{w}_{\text{new}})$
20:     Reset GNN parameters: $\theta_{\text{gnn}}\leftarrow\theta_{\text{gnn}}^{(0)}$
21:end procedure
22:procedure EditModel($\{(s,r,o)\},\mathbf{w}$)
23:     Input: Triples $\{(s,r,o)\}$, model weights $\mathbf{w}$
24:     Output: Updated weights $\mathbf{w}_{\text{final}}$
25:     Construct hyperbolic graph (Line 1)
26:     do
[…]
28:         Update weights using Möbius method (Line 11)
29:         Stabilize and reset GNN (Line 16)
30:     while Editing criteria not met
31:end procedure